# Small Stickers, Big Meanings: A Multilingual Sticker Semantic Understanding Dataset with a Gamified Approach

### v1

@@ 8. Conclusion
In this work, we propose StickerQueries, a bilingual benchmark designed to support and evaluate sticker understanding and retrieval in realistic conversational scenarios. By leveraging a gamified annotation process, we build a dataset that captures rich, high-quality queries reflecting the subtle and often culturally grounded semantics of sticker use. Our experiments show that fine-tuning vision-language models on this dataset leads to substantial performance improvements in both query generation and retrieval tasks. These findings highlight the limitations of general-purpose models and the need for specialized datasets in high-context, multimodal applications. We believe StickerQueries provides a valuable foundation for future work in personalized visual communication, semantic understanding, and cross-lingual sticker interaction.

### v2

@@ 8. Conclusion
In this work, we propose StickerQueries, a multilingual sticker semantics understanding dataset for sticker understanding and retrieval, built alongside Sticktionary, a gamified annotation framework. Together, they capture rich, high-quality queries and culturally grounded sticker semantics. Fine-tuning vision-language models on our dataset improves query generation and retrieval, highlighting the limits of general-purpose models and the need for specialized data in high-context multimodal tasks. These resources provide a strong foundation for future work in personalized visual communication, semantic understanding, and cross-lingual sticker interaction.